# Extrapolation of Machine-Learning Interatomic Potentials for Organic and Polymeric Systems

### v2

@@ Abstract
Machine-Learning Interatomic Potentials (MLIPs) have surged in popularity due to their promise of expanding the spatiotemporal scales possible for simulating molecules with high fidelity. The accuracy of any MLIP is dependent on the data used for its training; thus, for large molecules, like polymers, where accurate training data is prohibitively difficult to obtain, it becomes necessary to pursue non-traditional methods to construct MLIPs, many of which are based on constructing MLIPs using smaller, analogous chemical systems. However, we have yet to understand the limits to which smaller molecules can be used as a proxy for extrapolating macromolecular energetics. Here, we provide a “control study” for such experiments, exploring the ability of MLIP approaches to extrapolate between n=1-8 n-polyalkanes at identical conditions. Through Principal Covariates Classification, we quantitatively demonstrate how convergence in chemical environments between training and testing datasets coincides with an MLIP’s transferability. Additionally, we show how careful attention to the construction of an MLIP’s neighbor list can promote greater transferability when considering various levels of the energetic hierarchy. Our results establish a roadmap for how one can create transferable MLIPs for macromolecular systems without the prohibitive cost of constructing system-specific training data.
I Introduction
Figure 1: How well do ML potentials built on short-chain molecules generalize to their longer counterparts?
In molecular dynamics, the definition of “potential energy” is one of the most critical simulation components, determining both the relevance and performance of the resulting computational experiment. Thus, there are countless methods for determining the potential energy of a chemical system; these methods often play the balance between assumed accuracy and efficiency, with the more “accurate” methods focused on calculating or approximating the underlying electronic structure of a system, whereas the more “efficient” methods focus on simple mathematical functions that can be computed easily for many atoms across long length and time scales.
For polymeric systems, which typically contain molecules consisting of hundreds or thousands of atoms, the latter, more “efficient” approach, classical force fields, has been the necessary choice. In a classical force field, we assume that a potential energy function can be broken down into conceptually discrete components corresponding to bonds, torsions, dihedrals, and van der Waals interactions. This results in a simple, fixed-form potential energy function parameterized to fit a specific system, for which there are countless options[1], approaches, and strategies, to varying degrees of experimental fidelity and success. While these force-fields aim to be accurate, many struggle to generalize beyond their original reference system. Furthermore, they often require the use of specific “scale” or “fudge” factors to improve experimental agreement[2, 3]. As such, it is often hard to trust classical force fields “as-is,” and thus the results obtained from them are not considered representative until experimentally verified.
Recent years has seen an explosion of “machine-learned interatomic potentials” (MLIPs), which recast the “accurate” into “efficient” procedures, providing surrogate models to electronic structure calculations that can be computed on-the-fly[4, 5, 6, 7, 8, 9, 10, 11]. Such MLIPs not only provide accurate and fast energetic predictions, but several strategies have been utilized that improve the generalizability of these models[12, 13, 14]. Similar to traditional force-field fitting, MLIPs work by determining potential energy functions for molecular configurations from a set of training configurations. The power in these approaches is their flexibility in the potential energy functional form, alleviating the source of many errors in classical force fields and enabling the study of ionic transport[15, 16], condensed-phase phenomena[17, 18], and electronic properties[19].
A challenge in applying these approaches to macromolecular systems stems from the fact that MLIPs, particularly those built for a specific system, require high-quality, ab initio training data on which to build a machine-learning model. This, in itself, is prohibitive for macromolecular systems, where such calculations are less feasible. Thus, despite the many problems in polymer science that require a quantum chemical level of accuracy, specially-trained MLIPs are out of reach[20].
So, how do we harness these powerful methods for polymeric simulation, where it is difficult, if not infeasible, to construct the requisite underlying training data? “Universal” or “foundational” MLIPs (UMLIPs) have recently emerged as a solution, as they promise to yield accurate potential energies for a wide variety of systems from pre-trained models[21, 22, 23, 24]. These potentials, owing to their recency, have yet to be tested for polymeric systems. Yet, while UMLIPs are generally applicable to many systems, there will always exist cases where custom training will be required.
To create a custom MLIP for polymeric systems, a popular approach is molecular fragmentation[25], in which one performs ab-initio calculations on small molecular fragments and combines them to calculate the energy of larger molecules. The molecular fragmentation approaches have been utilized within MLIP frameworks such as QRNN[26, 27] and GEMS[28]. These combinations yielded a smooth PES that resulted in stable and accurate molecular dynamics simulations. However, while powerful, this approach still requires explicit classical simulation of the overarching system to generate fragments, for which FFs are prone to their standard pitfalls.
As many macromolecular molecules include consistent, repeated chemistries, we hypothesize that we can create an MLIP trained on smaller analogs of a macromolecule and extrapolate the energetics to longer systems, as posited in Fig. 1. To test this hypothesis, we curate a dataset specifically for this purpose – n=18n=1-8 linear alkanes simulated at similar conditions (i.e., phase, density, pressure, and temperature). By doing so, we can provide critical guidance to MLIP construction for polymeric systems, both using traditional approaches and molecular fragmentation.
Our analysis shows that extrapolation is possible once all relevant carbon and hydrogen environment statistics have converged within the training set. This result, while statistically demonstrated, is nuanced, as extrapolating pre-built potentials requires additional procedures beyond those typically practiced in the community. And thus, we end our discussion with practical steps how to extrapolate polymeric MLIPs.
II Methods
II.1 Data Generation
@@
Each MLIP considered was trained on a single n=18n=1-8 $n$-polyalkane dataset. To limit confounding factors in our analyses, datasets were constructed at similar phase, pressure, and temperature conditions. Each dataset consisted of $N=64$ molecules, simulated at T=300T=300K with the volume corresponding to NIST fluid properties’ density for each alkane at 5 MPa and 300K[29]. At these temperature and pressure conditions, each alkane is in a liquid state, aside from methane, which is a supercritical liquid. Initial configurations of n=38n=3-8 polyalkanes were obtained from the Cambridge Structural Database[30]. For methane and ethane, initial configurations were generated using RDKit[31] on a cubic lattice with an appropriate volume according to NIST fluid properties’ data. Each initial configuration was relaxed to an unordered, liquid state using isobaric isothermal (NPT) molecular dynamics simulations at low pressure ($P=0.5$ MPa, T=300T=300K ) using density-functional tight binding (DFTB+)[32] and the mio Slater-Koster files[33]. These NPT simulations were performed until the appropriate density corresponding to NIST fluid properties’ data was obtained for each polyalkane. For $n=7,8$ datasets, further relaxation at large box volume (with $P=0.01$ MPa, T=300T=300K) was necessary in order to remove initialization artifacts. After these systems were randomized, LAMMPS[34] was used to repressurize these two datasets to 300K and 5 MPa with the OPLS-AA forcefield until they reached the appropriate volume [35]. With $n=8$, some initialization artifacts still remained, so we computed structure optimization in ASE using a BFGS algorithm and fmax=0.05 with a DFTB+ calculator using a Conjugated Gradient driver to anneal the system.
@@
NVT simulations were then conducted at 300K. These production simulations were conducted using DFTB+, initially using a grid of (1, 1, 1) k-points, and were run until potential energies had converged. For our purposes, an equilibrium cutoff was defined as where the running average of the total energy did not fluctuate by more than 0.1 eV for 2,000 consecutive frames in the NVT simulation. We then selected 1000 frames from the subsequent simulation at random to form testing sets, and used farthest-point sampling[36, 37, 38] on Smooth Overlap of Atomic Positions (SOAP[4]) vectors to select a mutually-exclusive 10,000 frames to constitute our training sets. We recomputed the energies and forces for the training and testing datasets using higher grid-point interpolation (3, 3, 3) using a high-throughput workflow with signac[39] and signac-flow[40].
@@
II.2 MLIP Training
For every MLIP, we solely used the aforementioned “training” sets during construction, and trained each model as one normally would for a non-extrapolative task. When training solely on intramolecular quantities (Fig. S3), we down-selected from the 640,000 single-molecule frames to 10,000, choosing one random molecule from each training frame. Testing sets were kept consistent, with 64,000 single molecule frames in total.
II.2.1 MACE
The MACE[5] model (Higher‑Order Equivariant Message Passing Neural Network) combines equivariant graph neural network design with explicit many‑body message passing. Unlike most standard MPNNs that rely on only two‑body (pairwise) interactions and require many message‑passing layers to build complexity, MACE constructs higher‑body (up to four‑body) equivariant messages in a hierarchical fashion. This allows the network to achieve high representational power with just two message‑passing iterations, substantially reducing depth, improving parallelism, and maintaining scalability while reaching or surpassing state‑of‑the‑art accuracy on benchmarks such as rMD17, 3BPA, and AcAc[5]. By building hierarchical tensor products of lower‑order features $A^{(t)}_{i}$ and symmetrizing via generalized Clebsch–Gordan coupling, MACE efficiently implements these many‑body messages without exponential cost. This induces equivariant feature tensors $B^{(t)}_{i,\eta\nu,kLM}$, which are linearly combined (with weights dependent on element type and irreducible representation indices $k,L,M$) to produce message updates $m^{(t)}_{i,kLM}$. For our MACE-based potentials, we use the software v.0.3.13 and a standard set of model hyperparameters, as shown in the example script in App. A.
(a) Energy-fitting results of MACE MLIP for different train (x-axis) and test (y-axis) pairings, respectively. Color indicates the mean-absolute-error (MAE) in units of meV/atom, respectively.
(b) Force-fitting results of MACE MLIPs for different train (x-axis) and test (y-axis) pairings, respectively. Color indicates the mean-absolute-error (MAE) in units of meV/atom, respectively.
@@ II.2.1 MACE
II.2.2 SOAP Analysis
Smooth Overlap of Atomic Positions (SOAP) vectors were computed using the featomic[41] and librascal [42] packages with an interaction cutoff of $5\AA$ (for Fig. 3, Fig. S4) and $7\AA$ (for Fig. 4, Fig. S5); the GTO basis set was computed with 8 radial components and 4 angular components. Principal Covariates Classification (PCovC, [43]) was computed using v0.3.2 of scikit-matter[36], using as $X$ the aforementioned SOAP vectors of the training set, a mixing parameter of $0.5$, a logistic regression classifier, and oligomer length $n$ as our class value. Each PCovC model was computed solely for like-environments, such as H-, CH2, or CH3 environments (Fig. 3, Fig. S4). Prior to fitting and analysis, feature vectors and energies were appropriately scaled and centered using scikit-matter[36] StandardFlexibleScaler. For regression (Fig. 4), we used scikit-learn[44] RidgeCV with cv=5 and a negative mean squared error scorer. 20 regularization constants ($\alpha$) were considered for RidgeCV using a logarithmic grid.
III Results
On first glance, the energy-fitting results of Fig. 2(a) tell a discouraging story – each training set is unable to predict the energies of the other polyalkanes, with a MAE well-above chemical accuracy (which we define as 1meV/atom[45]). However, the force results (Fig. 2(b)) imply something much different, in line with our chemical intuition – at larger chain lengths, we can reasonably extrapolate the energetics of one system to the next (and even those below).
@@ III Results
However, this is not the case here, where the $E_{C}-E_{H}$ is on the order of 30eV (see App. A), leading to a much larger value than the observed shift. This presents an issue for extrapolating energetics from one system to another, where the only information known a priori may be the difference in underlying chemistry, and where “transfer learning” may be infeasible due to the lack of underlying training data. Thus, we first determine if $b$ is a linear function of our composition ($x_{C}$, $x_{H}$), as this would still allow for the extrapolation of one system to another without prior computation of energetics or configurations.
@@
For models trained on n=48n=4-8 alkanes, there exists a strong linear proportionality (Pearson correlation coefficient = 0.9997, Fig. 2(c)) between the compositional shift and the mean-shift of the datasets. This suggests that these mean-shifts are learnable parameters, even if not obviously grounded in physical values such as $E_{atom}$. To test this, we determined a slope to the line in Fig. 2(c) via non-regularized linear regression, omitting from training the values obtained from models built on methane (blue), ethane (light blue), and propane (green), which are considered ill-conditioned models. Accounting for the large co-linearity of the mean-shift and compositional change, many of the error predictions are improved, yet there exist some residual errors (Fig. 2(d)). This suggests that the change in conformational distribution contributes non-negligibly to the mismatch in energies.
@@
However, the results of Fig. 2(b) still demonstrate that there exist two critical lengths of polyalkane for extrapolation: 1) butane, where we see a general reduction in error magnitude from 20-30meV/Å to 3-6meV/Å (omitting methane), and 2) hexane, after which we see little improvement in force error with additional molecular units. The change in errors from propane to butane follows our chemical intuition – prior to this chain length, there is no sampling of the dihedral rotation possible in these molecules, and thus any system (n=48n=4-8) that contains these interactions would be ill-represented by such models.
But what makes hexane special? To understand this, we employ hybrid supervised-unsupervised analyses on the underlying environments within the training sets, focusing on the three-body interactions of the CH2 terms, which demonstrate the greatest statistical difference from dataset to dataset (see Fig. S4 for further analyses). We do so by employing a preceding technology to MACE descriptors, the Smooth Overlap of Atomic Positions (SOAP[4]), which are deterministic in their description of atomic environments (as opposed to MACE descriptors, whose higher-order correlations are determined through the message-passing component of the MACE training). We consider only the descriptors based on the first layer of the MACE perceptive field (5.0Å) that will be most heavily weighted in the constructed regression. We pass the SOAP vectors through Principal Covariates Classification (PCovC)[43, 46, 36], which computes the two-dimensional projection that both retains PCA-like variance and best separates the data based on assigned labels. In this context, PCovC can be applied to identify the environments that distinguish one training set from another (by using the training set identities as our classification labels). As is shown in Fig.  S4, a PCA projection is unable to yield such delineation in similar datasets.
Figure 3: Distribution of CH2 environments for the different training sets, mapped Principal Covariates Classification (PCovC[43]). The map demonstrates the convergence of environments with increasing $n$.
@@
From this analysis, we can see that the convergence in MLIP error corresponds with the sampling of “E” environments with hexane, and the addition of “F” environments within heptane and octane leads to diminishing returns in terms of force approximation. This demonstrates that determining the minimal chain length required for MLIP construction (or molecular fragmentation[25, 26, 27, 28]) amounts to determining the convergence of the majority of molecular environments, a task that can be accomplished via classification models or similar unsupervised tasks on single-molecule data.
@@
However, predicting the total energies and forces may not be sufficient for a usable potential model, as the hierarchical nature of interactions within these systems decides much of their fundamental behavior[13]. Here, we consider potential energies emerging from three distinct classes of interactions:
@@
MACE itself (v.0.3.13) is not natively suited for directly learning these interactions, nor, would we argue, are atom-centered expansions that implicitly or explicitly focus on the nearby atoms in constructing their MLIP. To demonstrate this, we again turn to the precursor for MACE[5], SOAP[4], and show how careful attention to weighting interactions can yield better fidelity for intermolecular models.
Following the same procedures in terms of training/testing splits, we construct cross-validated Ridge Regressions on SOAP representations for the same n=18n=1-8 $n$-polyalkane datasets used for MACE MLIP construction. SOAP vectors were calculated with a cutoff of 7Å for atomic environments and averaged for each frame such that for the 10,000 training frames and 1,000 testing frames, we have 10,000 and 1,000 corresponding SOAP vectors; we refer to these initial SOAP calculations as the “total” SOAP vectors, or $X^{\text{total}}$.
@@
Similar to the initial results from MACE in Fig. 2(a), the results of using a $E_{\text{intermolecular}}\approx X^{\text{total}}w$, shown in Fig. 4(a), are discouraging. Extrapolation, with a threshold MAE of 1 meV/atom [45], is only achieved three times (by hexane on heptane, and heptane on hexane/octane). However, we can adopt lessons from learning the lattice energy of molecular crystals[48] – SOAP representations, and representations that are implicitly or explicitly “near-sighted”[5, 49, 50] will, by design, consider the contributions of nearby atoms more strongly than those further away. This will overshadow the intermolecular environments which contribute to $E_{intermolecular}$, limiting our accuracy in extrapolation.

### v3

@@ Abstract
Machine-Learning Interatomic Potentials (MLIPs) have surged in popularity due to their promise of expanding the spatiotemporal scales possible for simulating molecules with high fidelity. The accuracy of any MLIP is dependent on the data used for its training; thus, for large molecules like polymers and biomolecules, where ab initio training data is prohibitively difficult to obtain, it becomes necessary to use smaller, analogous chemical systems to construct MLIPs. Here, we perform a control study using n=18n=1-8 linear alkanes to determine when MLIPs trained on small molecules can accurately extrapolate to longer chains and more complex architectures. By combining MLIP performance analysis with environment‑resolved SOAP descriptors and Principal Covariates Classification, we show that reliable extrapolation emerges precisely when the relevant local chemical environments have converged between training and target systems. We further demonstrate that careful construction of the neighbor list substantially improves the learnability of intermolecular energetics, the component most critical for polymeric behavior. Together, these results provide a practical, data‑driven blueprint for designing transferable MLIPs for macromolecular materials, whether built from bespoke training data or adapted from emerging universal MLIP frameworks.
Introduction
In molecular simulation, the choice of potential energy model determines both the accuracy of predicted behavior and the computational time scales that can be achieved. Classical force fields have played a foundational role in polymer science for decades, enabling the study of thermodynamics, structure, and dynamics across a wide variety of molecular and polymeric systems. Their conceptual clarity, expressing molecular energetics through well‑defined bonded and nonbonded terms, combined with their computational efficiency, continues to make them indispensable for simulations that require long time scales, large system sizes, or broad chemical screening.
At the same time, advances in electronic‑structure theory, machine learning, and high‑performance computing have created new opportunities at the interface of accuracy and efficiency. Machine‑learned interatomic potentials (MLIPs) offer one such opportunity: by learning from quantum‑chemical reference data, they can capture complex many‑body interactions without requiring an explicit decomposition into fixed functional forms[2, 4, 6, 11, 5, 38, 36, 37]. MLIPs have recently demonstrated strong performance across problems ranging from ionic transport[49, 19], condensed-phase phenomena[10, 52], and electronic properties[23], motivating interest in when and how these models can complement existing force‑field approaches–particularly for systems where non-additive electronic effects or chemical heterogeneity play an important role.
However, using MLIPs for large molecules such as polymers or biomolecules remains challenging. High‑quality quantum chemical data are difficult to obtain for such macromolecules[50], and transferability across chain length or architecture is not guaranteed or well-understood. This has led to increasing interest in strategies where smaller, chemically related molecules act as surrogates for constructing potentials that are then applied to larger targets[29, 41, 48]. Although intuitive, the limits of such extrapolation have not been systematically quantified.
Universal or “foundational” machine-learned interatomic potentials (UMLIPs) have recently emerged as a promising direction, offering broad chemical coverage from large, diverse training sets[40, 3, 45, 51]. These models reduce the need for system‑specific quantum data and have shown strong performance across many small‑molecule and materials benchmarks. However, due to their recency, their behavior for macromolecular systems remains largely untested. Moreover, even when UMLIPs provide a reliable baseline, fine‑tuning or system‑specific adaptation is often needed for application-specific studies. And thus, understanding which molecular environments must be represented is still quintessential, regardless of whether one uses a bespoke MLIP or a UMLIP.
In this work, we provide a control study of this question. Using linear alkanes with n=18n=1-8 as a model system, we examine when MLIPs trained exclusively on small molecules can accurately extrapolate to longer chains and more complex architectures. By combining MLIP performance analysis with environment‑level descriptors and classification tools, we identify the molecular environments that govern extrapolation accuracy and show how neighbor‑list design influences the learnability of inter‑ versus intramolecular interactions. These results form a practical blueprint for building transferable MLIPs for macromolecular systems, in a way that complements, and in some cases extends, the capabilities of traditional force‑field approaches.
Methods
Data generation
Training sets
@@ Methods
Each MLIP was trained on a single n=18n=1-8 $n$-polyalkane dataset. To limit confounding factors in our analyses, datasets were constructed at similar phase, pressure, and temperature conditions. Each dataset consisted of $N=64$ molecules, simulated at T=300T=300K with the volume corresponding to NIST fluid properties’ density for each alkane at 5 MPa and 300K[39]. At these temperature and pressure conditions, each alkane is in a liquid state, aside from methane, which is a supercritical liquid. Initial configurations of n=38n=3-8 polyalkanes were obtained from the Cambridge Structural Database[24]. For methane and ethane, initial configurations were generated using RDKit[35] on a cubic lattice with the appropriate volume. Each initial configuration was relaxed to an unordered, liquid state using isobaric-isothermal (NPT) molecular dynamics simulations at low pressure ($P=0.5$ MPa, T=300T=300K). These NPT simulations were performed with a Berendsen barostat and thermostat, a timestep of 0.2 femtoseconds as well as periodic boundary conditions, and were run until the reference experimental density was obtained for each polyalkane.
For $n=7,8$ datasets, further relaxation at large box volume (with $P=0.01$ MPa, T=300T=300K) and geometry optimization was necessary in order to remove initialization artifacts. After these systems were randomized, we repressurized these two datasets to 300K and 5 MPa with the OPLS-AA forcefield until they reached the appropriate volume [32]using periodic boundary conditions, a timestep of 1 femtosecond, and a Nose-Hoover thermostat and barostat. Initial velocities were generated using a Maxwell-Boltzmann sampling at 300 K without fixing a random seed. With n = 8, some initialization artifacts still remained, so we computed structure optimization in ASE using a BFGS algorithm and fmax=0.05 with a DFTB+ calculator using a Conjugated Gradient driver to anneal the system.
@@
After the appropriate box volumes were reached, NVT simulations were then conducted at 300K. These production simulations were conducted using DFTB+, initially using a grid of (1, 1, 1) k-points with a Berendsen thermostat, an initial stochastic Maxwell-Boltzmann velocity distribution, a timestep of 0.2 femtoseconds, as well as periodic boundary conditions, and were run until potential energies had converged. For our purposes, an equilibrium cutoff was defined as where the running average of the total energy did not fluctuate by more than 0.1 eV for 2,000 consecutive frames in the NVT simulation. We then selected 1000 frames from the subsequent simulation at random to form testing sets, and used farthest-point sampling[21, 28, 8] on Smooth Overlap of Atomic Positions (SOAP[2]) vectors to select a mutually-exclusive 10,000 frames to constitute our training sets.
Additional testing sets
Decane, dodecane, cyclohexane, 4-propylheptane, and 3,3-diethylpentane, configurations were sampled using the OPLS all-atom (AA) forcefield[32]. Decane and dodecane were simulated with 48 molecules, and the others with 64 molecules, in order to keep computational overhead consistent. The molecules were initialized on a lattice with 10 Å spacing and with each dimension of each lattice position randomly perturbed by -5 to 5 Å sampled from a uniform distribution. We equilibrated the systems with respect to the average radius of gyration, which quantifies the average size of the alkane conformations. The radius of gyration of an alkane, $R_{g}$, is given by $R_{g}=\sqrt{\frac{\sum_{i}m_{i}r_{i}^{2}}{\sum_{i}m_{i}}},$ where the sums are over atoms $i$ in the alkane, $m_{i}$ is the atomic mass, and $r_{i}$ is the displacement relative to the alkane center of mass. We used pyMBAR [12] to calculate the correlation time of the radius of gyration time series. We equilibrated each system for 100 correlation times. Figure S6 demonstrates the equilibrium radius of gyration of the alkanes. Once equilibrated, the systems were run for 1 ns. All classical simulations were run in LAMMPS [47] using forcefield parameters generated by ligpargen[33, 14, 13] in an NPT ensemble at 300 K and 5 MPa using a Nose-Hoover thermostat and barostat with loose couplings of 1 $\text{ps}^{-1}$ each and a timestep of 1 fs.
Energy and force evaluations
For all frames in the training and testing data, we recomputed the energies and forces with DFTB+ using self-consistent charges and the mio Slater-Koster files[17] using a Monkhurst-Pack grid-point interpolation (3, 3, 3). We utilized a high-throughput workflow with signac[1] and signac-flow[46].
@@
Machine-learned potentials and analyses
MACE MLIPs
For every MLIP, we solely used the aforementioned training sets during construction, and trained each model as one normally would for a non-extrapolative task. When training solely on intramolecular quantities (Fig. S3), we down-selected from the 640,000 single-molecule frames to 10,000, choosing one random molecule from each training frame. Testing sets were kept consistent, with 64,000 single molecule frames in total.
The MACE[4] model (Higher‑Order Equivariant Message Passing Neural Network) combines equivariant graph neural network design with explicit many‑body message passing. Unlike most standard MPNNs that rely on only two‑body (pairwise) interactions and require many message‑passing layers to build complexity, MACE constructs higher‑body (up to four‑body) equivariant messages in a hierarchical fashion. This allows the network to achieve high representational power with just two message‑passing iterations, substantially reducing depth, improving parallelism, and maintaining scalability while reaching or surpassing state‑of‑the‑art accuracy on benchmarks such as rMD17, 3BPA, and AcAc[4]. By building hierarchical tensor products of lower‑order features and symmetrizing via generalized Clebsch–Gordan coupling, MACE efficiently implements these many‑body messages without exponential cost. This induces equivariant feature tensors, which are linearly combined (with weights dependent on element type and irreducible representation indices to produce message updates. For our MACE-based potentials, we use the software v.0.3.13 and a standard set of model hyperparameters, as shown in the example script in the Appendix.
(a) Results of predicting per-configuration energies (normalized by the number of atoms) using MACE MLIPs for different train (x-axis) and test (y-axis) pairings, respectively. Color indicates the mean-absolute-error (MAE) in units of meV/atom, respectively.
(b) Results of predicting atomic force-vectors using the MACE MLIPs for different train (x-axis) and test (y-axis) pairings, respectively. Errors correspond to the average magnitude of vectorial error across the testing set. Color indicates the mean-absolute-error (MAE) in units of meV/Å, respectively.
@@
SOAP Analysis
Smooth Overlap of Atomic Positions (SOAP) vectors were computed using the featomic[7] and librascal [42] packages with an interaction cutoff of 1010Å (for Fig. 2, Fig. S4) and 77Å (for Fig. 3, Fig. S5). In all cases, we used the GTO basis set with 8 radial components and 4 angular components. Principal Covariates Classification (PCovC, [30]) was computed using v0.3.2 of scikit-matter[21], using as $X$ the aforementioned SOAP vectors of the training set, a mixing parameter of $0.5$, a logistic regression classifier, and oligomer length $n$ as our class value. Each PCovC model was computed solely for like-environments, such as hydrogen, CH2, or tertiary carbon environments (Fig. 2, Fig. S4). Prior to fitting and analysis, feature vectors and energies were appropriately scaled and centered using scikit-matter[21] StandardFlexibleScaler. For regression (Fig. 3), we used scikit-learn[44] RidgeCV with cv=5 and a negative mean squared error scorer. 20 regularization constants ($\alpha$) were considered for RidgeCV using a logarithmic grid.
Results
Extrapolating energies and forces
Fig. 1 shows the results of using trained MLIPs to predict the energies and forces on different oligomeric chain lengths. On first glance, the energy-fitting results of Fig. 1(a) tell a discouraging story – each training set is unable to predict the energies of the other polyalkanes, with a MAE well-above chemical accuracy (which we define as 1meV/atom[53]). However, the force results (Fig. 1(b)) imply something much different, in line with our chemical intuition – at larger chain lengths, we can reasonably extrapolate the energetics of one system to the next (and even those below).
@@ Results
However, this is not the case here, where a shift due solely to atomic contributions would lead to a much larger value than the observed shift. This presents an issue for extrapolating energetics from one system to another, where the only information known a priori may be the difference in underlying chemistry, and where “transfer learning” may be infeasible due to the lack of underlying training data. Thus, we first determine if $b$ is a linear function of our composition ($x_{C}$, $x_{H}$), as this would still allow for the extrapolation of one system to another without prior computation of energetics or configurations.
@@
For models trained on n=48n=4-8 alkanes, there exists a strong linear proportionality (Pearson correlation coefficient = 0.9997, Fig. 1(c)) between the compositional shift and the mean-shift of the datasets. This suggests that these mean-shifts are learnable parameters, even if not obviously grounded in physical values such as $E_{atom}$. To test this, we determined a slope to the line in Fig. 1(c) via non-regularized linear regression, omitting from training the values obtained from models built on methane (blue), ethane (light blue), and propane (green), which are considered unsuccessful models. Accounting for the large co-linearity of the mean-shift and compositional change, many of the error predictions are improved, yet there exist some residual errors (Fig. 1(d)). This suggests that the change in conformational distribution contributes non-negligibly to the mismatch in energies.
@@
However, the results of Fig. 1(b) still demonstrate that there exist two critical lengths of polyalkane for extrapolation: 1) butane, where we see a general reduction in force error magnitude from 20-30meV/Å to 3-6meV/Å (omitting methane), and 2) hexane, after which we see little improvement in force error with additional molecular units. The change in errors from propane to butane follows our chemical intuition – prior to this chain length, there is no sampling of the dihedral rotation possible in these molecules, and thus any system (n=48n=4-8) that contains these interactions would be ill-represented by such models.
But what makes hexane special? To understand this, we can look at the higher-dimensional representation of the interaction environments shared and distinct within the training sets. While we could arguably use the learned features of the MACE MLIPs, we instead employ a preceding, related technology, the Smooth Overlap of Atomic Positions (SOAP[2]), which are deterministic in their description of atomic environments without the implications of training on one dataset versus another . Both MACE and SOAP vectors compute the many-body expansion of atomic interactions, and thus we can infer some behaviors of MACE based on a SOAP representation. We construct our SOAP descriptors based on the distance of the MACE perceptive field (10.0Å).
After we compute the SOAP vectors for each training set, a standard analytical task is to look at the Principal Components Analysis (PCA) of all vectors, which would demonstrate the diversity of environments. However, here we are particularly interested in how the environments within these datasets differ, which is not the underlying mathematical motivation of PCA, and this aspect can be largely lost in PCA (and other unsupervised) maps. So, we instead employ hybrid supervised-unsupervised analyses called Principal Covariates Classification (PCovC)[30, 26, 22, 21], which computes the two-dimensional projection that both retains PCA-like variance and best separates the data based on assigned labels (here the number of carbons in the dataset backbone). Here we focus on the three-body interactions of the CH2 terms, which demonstrate the greatest statistical difference from dataset to dataset, with PCA and PCovC maps of remaining environments in Fig. S4.
Figure 2: Distribution of CH2 environments for the different training sets, mapped via Principal Covariates Classification (PCovC[30]). The axes are directions in latent space that demonstrate the highest degree of diversity in the CH2 environments and separability between the datasets. Points that are overlapping are considered “indistinguishable,” in that an ML model could not tell whether the environment sits in one alkane length or another. The map demonstrates the convergence of environments with increasing $n$. Different alkanes are denoted in different colors (green - propane, yellow - butane, orange - pentane, red - hexane, purple - heptane, and brown - octane), with the corresponding CH2 environment enlarged for each cluster in the map.
@@
From this analysis, we see that the convergence in MLIP error corresponds with the sampling of the CH2 environments three carbons from the end, and the addition of further internal environments within heptane and octane leads to diminishing returns in terms of force approximation. This demonstrates that determining the minimal chain length required for MLIP construction (or molecular fragmentation[20, 29, 41, 48]) amounts to determining the convergence of the majority of molecular environments, a task that can be accomplished via classification models or similar unsupervised tasks on single-molecule data.
Separating the different contributions to total energies
@@
However, predicting the total energies and forces may not be sufficient for a usable potential model, as the hierarchical nature of interactions within these systems decides much of their fundamental behavior[15]. Conceptually, and in classical force fields,, we consider potential energies emerging from three distinct classes of interactions:
@@
MACE itself (v.0.3.13) is not natively suited for directly learning these interactions, nor are atom-centered expansions that implicitly or explicitly focus on the nearby atoms in constructing their MLIP[9]. To demonstrate this, we again turn to the SOAP[2] framework, and show how careful attention to weighting interactions can yield better fidelity for intermolecular models. Following the same procedures in terms of training/testing splits, we construct cross-validated Ridge Regressions on SOAP representations for the same n=18n=1-8 $n$-polyalkane datasets used for MACE MLIP construction. SOAP vectors were calculated with a cutoff of 7Å for atomic environments and averaged for each frame such that for the 10,000 training frames and 1,000 testing frames, we have 10,000 and 1,000 corresponding SOAP vectors; we refer to these initial SOAP calculations as the “total” SOAP vectors, or $X^{\text{total}}$.
@@
Similar to the initial results from MACE in Fig. 1(a), the results of using a $E_{\text{intermolecular}}\approx X^{\text{total}}w$, shown in Fig. 3(a), are discouraging. Extrapolation, with a field-standard threshold MAE of 1 meV/atom [53], is only achieved three times (by hexane on heptane, and heptane on hexane/octane), even despite the considerably smaller magnitude of the intermolecular energetics, as compared to the total energies regressed earlier. However, we can adopt lessons from learning the lattice energy of molecular crystals[9] – SOAP representations, and representations that are implicitly or explicitly “near-sighted”[4, 16, 25] will, by design, consider the contributions of nearby atoms more strongly than those further away. This will overshadow the intermolecular environments which contribute to $E_{intermolecular}$, limiting our accuracy in extrapolation.
(a) Force errors resulting from testing n=1-8 trained MLIPs on longer chain lengths and architectures. Color of the bar plot denotes the training set used, and each set of bars refers to a different test molecule.
(b) Distribution of force errors for cyclohexane (left), 4-propylheptane(center), and 3,3-diethylpentane (right), as predicted by a model trained on octane. Colors indicate the atoms within the molecule. For every carbon, the corresponding hydrogens are in the same hue, with a lighter tint.
(c) Intermolecular energy errors resulting from testing n=1-8 trained SOAP-Ridge MLIPs on longer chain lengths and architectures. Color of the bar plot denotes the training set used, and each set of bars refers to a different test molecule. Hatched bars indicate MLIPs constructed using far-sighted SOAP vectors $X^{fs}$, solid bars indicate MLIPs using total SOAP vectors.
Figure 4: Force and energy errors for decane, dodecane, and the nonlinear alkanes.